1) Show that there are two points of intersection of the hyperbola, with equation y= -8/x and the line y= -2x + 6, and hence find them.
Answer by lwsshak3(11628)

Show that there are two points of intersection of the hyperbola, with equation y= -8/x and the line y= -2x + 6, and hence find them.
***
-8/x=-2x+6
-8=-2x^2+6x
2x^2-6x-8=0
x^2-3x-4=0
(x-4)(x+1)=0
x=4,-1
y=-8/x
y=-2,8
points of intersection: (4,-2) and (-1,8)